# A Survey on Diffusion Models for Time Series and Spatio-Temporal Data

### v3

@@ 2.3 Typical Diffusion Models
Typically, the training process includes two steps: the forward process (diffusion) and the reverse process (denoising). Diffusion models start with a noise distribution, which is gradually altered through a series of steps. In the forward process, the model incrementally adds noise to the original data over multiple steps until the data turns into pure random noise. This process is usually Markovian, meaning that each step depends only on the preceding one. Then the reverse process takes place, involving learning to remove the noise from the data, essentially reversing the forward process. By training the model to remove the noise added during the diffusion process, the model learns to generate samples from the same distribution as the training data. The entire training process involves optimizing the model to denoise effectively. This is typically done using a loss function that encourages the model to produce samples that are close to the true data distribution [3].
@@
The currently common frameworks for diffusion models include denoised diffusion probabilistic models (DDPMs) [1, 2], score-based stochastic differential equations (Score SDEs) [6, 4], conditional diffusion models [7, 8, 9], etc. Following, we will introduce the subclasses of the diffusion model through theoretical formula derivation.
2.3.1 Denoised Diffusion Probabilistic Models (DDPMs)
@@
$q(\mathbf{x}_{k}|\mathbf{x}_{k-1}):=\mathcal{N}(\mathbf{x}_{k};\sqrt{1-\beta_{k}}\mathbf{x}_{k-1},\beta_{k}\mathbf{I}),$ (1)
@@
$q(\mathbf{x}_{k}|\mathbf{x}_{0}):=\prod_{k=1}^{K}{q}(\mathbf{x}_{k}|\mathbf{x}_{k-1}):=\mathcal{N}(\mathbf{x}_{k};\sqrt{\bar{\alpha_{k}}}\mathbf{x}_{0},\sqrt{1-\bar{\alpha_{k}}}\mathbf{I}),$ (2)
$\mathrm{where}\hskip 2.84526pt\alpha_{k}:=1-\beta_{k},\hskip 2.84526pt\mathrm{and}\hskip 2.84526pt\bar{\alpha_{k}}:=\prod_{i=1}^{K}\alpha_{i}.$ (3)
Therefore, $\mathbf{x}_{k}=\sqrt{\bar{\alpha_{k}}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha_{k}}}\epsilon$ with Gaussian noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. Typically, it is designed $\bar{\alpha_{k}}\approx 0$, s.t., $q(\mathbf{x}_{k}):=\int q(\mathbf{x}_{K}|\mathbf{x}_{0})q(\mathbf{x}_{0})\mathbf{d}\mathbf{x}_{0}\approx\mathcal{N}(\mathbf{x}_{k};\mathbf{0},\mathbf{I})$, i.e., the backward chain can begin with any Gaussian noise. Overall, the forward process gradually injects noise into the data until all structures have disappeared.
@@
${p}_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k}):=\mathcal{N}(\mathbf{x}_{k-1};\mu_{\theta}(\mathbf{x}_{k},{k}),\sigma_{\theta}(\mathbf{x}_{k},{k})\mathbf{I}),$ (4)
@@
$\mathbb{E}\left[-\log p_{\theta}\left(\mathrm{x}_{0}\right)\right]$ $\leq\mathbb{E}_{q}\left[-\log\frac{p_{\theta}\left(\mathrm{x}_{0:K}\right)}{q\left(\mathrm{x}_{1:K}\mid\mathrm{x}_{0}\right)}\right]$ (5)
$=\mathbb{E}_{q}\left[-\log p\left(\mathrm{x}_{K}\right)-\sum_{k\geq 1}\log\frac{p_{\theta}\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k}\right)}{q\left(\mathrm{x}_{k}\mid\mathrm{x}_{k-1}\right)}\right]$
$=:L.$
It is equivalent to Kullback–Leibler divergence (KL divergence) format as mentioned in the DDPM paper [2] with three parts: the prior loss $L_{K}$, the divergence of the forwarding step and the corresponding reversing step $L_{k-1}$, and the reconstruction loss $L_{0}$:
$L:=$ $\mathbb{E}_{q}[\underbrace{D_{\mathrm{KL}}\left(q\left(\mathrm{x}_{K}\mid\mathrm{x}_{0}\right)\|p\left(\mathrm{x}_{K}\right)\right)}_{L_{K}}$ (6)
$+\sum_{k>1}\underbrace{D_{\mathrm{KL}}\left(q\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k},\mathrm{x}_{0}\right)\|p_{\theta}\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k}\right)\right)}_{L_{k-1}}$
$\underbrace{-\log p_{\theta}\left(\mathrm{x}_{0}\mid\mathrm{x}_{1}\right)}_{L_{0}}].$
@@
$q\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k},\mathrm{x}_{0}\right)=\mathcal{N}\left(\mathrm{x}_{k-1};\tilde{\boldsymbol{\mu}}_{k}\left(\mathrm{x}_{k},\mathrm{x}_{0}\right),\tilde{\beta}_{k}{I}\right),$ (7)
$\tilde{\mu}_{t}(\mathbf{x}_{k},\mathbf{x}_{0}):=\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\mathbf{x}_{0}+\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\mathbf{x}_{k}\hskip 2.84526pt,$ (8)
$\tilde{\beta}_{k}:=\frac{1-\bar{\alpha}_{k-1}}{1-\bar{\alpha}_{k}}\beta_{k}.$ (9)
@@
$L_{k-1}=\mathbb{E}_{q}\left[\frac{1}{2\sigma_{k}^{2}}\left\|\tilde{\boldsymbol{\mu}}_{k}\left(\mathrm{x}_{k},\mathrm{x}_{0}\right)-\boldsymbol{\mu}_{\theta}\left(\mathrm{x}_{k},k\right)\right\|^{2}\right]+C.$ (10)
Ho et al. [2] emphasize that, rather than parameterizing the mean $\mu_{\theta}(\mathrm{x}_{k},k)$, predicting the noise vector at each time step in the forward process by parameterizing $\epsilon_{\theta}(\mathrm{x}_{k},k)$ for simplification:
$\mathbb{E}_{k\sim\mathcal{U}(1,K),\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\bigg{[}{\lambda(k)\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{k},k)\right\|^{2}}\bigg{]},$ (11)
where $\lambda(k)=\frac{{\beta_{k}}^{2}}{2{\sigma_{K}}^{2}\alpha_{k}(1-\bar{\alpha}_{k})}$ is a weight that changes noise scale, and $\epsilon_{\theta}$ is a model for Gaussian noise prediction. After training using the above loss function, $\epsilon_{\theta}$ will be used in the reverse process of ancestor sampling.
@@
$p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k})$ $=\mathcal{N}(\mathbf{x}_{k-1};\mu_{\theta}(\mathbf{x}_{k},k),\sigma_{\theta}^{2}(\mathbf{x}_{k},k)\mathbf{I})$ (12)
$\sim\frac{1}{\sqrt{\alpha_{k}}}(\mathbf{x}_{k}-\frac{\beta_{k}}{\sqrt{1-\overline{\alpha_{k}}}}\epsilon_{\theta}(\mathbf{x}_{k},k))+\sigma_{\theta}(\mathbf{x}_{k},k)z,$
@@ 2.3.2 Score SDE Formulation
DDPM achieved a set of discrete steps in the forward processing, so it has some limitations about training designs. Score SDE further generalizes DDPM’s discrete system to a continuous framework based on the stochastical differential equation [6]. Here we use $T$ instead of the step size $k$ in DDPM.
@@
Forward Process. The corresponding continuous diffusion process can be represented using Itô SDE [291], including a mean shift and a Brownian motion (standard Wiener process) as follows:
$\mathrm{d}\mathbf{x}=\mathbf{f}(\mathbf{x},t)\mathrm{d}t+g(t)\mathrm{d}\mathbf{w},t\in[0,T],$ (13)
@@
where $\mathbf{f}(\cdot,t)$ represents the drift coefficient for the stochastic process $\mathbf{x}(t)$, and $g(\cdot)$ is the diffusion coefficient linked with the Brownian motion $\mathbf{w}$.
Similar with DDPM, $\mathbf{x_{0}}$ and $\mathbf{x}_{T}$ represent sequence from the clean distribution $p_{0}=\mathcal{N}(\mathbf{x}_{0};\mathbf{0},\mathbf{I})$ and the standard Gaussian distribution $p_{T}=\mathcal{N}(\mathbf{x}_{T};\mathbf{0},\mathbf{I})$, respectively. And the corresponding SDE is:
$\mathrm{d}\mathbf{x}=-\frac{1}{2}\beta(t){\mathbf{x}~{}d}t+\sqrt{\beta(t)}\mathrm{d}\mathbf{w}.$ (14)
@@
Reverse Process. The new samples can be synthesized from the known prior distribution $p_{T}$ by solving the reverse-time SDE [174]:
$\mathrm{d}\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]\mathrm{d}t+g(t)\mathrm{d}\bar{\mathbf{w}},$ (15)
@@
Where $\bar{\mathbf{w}}$ is a Brownian motion with reversed time flows [148]. The solution to the reverse-time SDE is approximated by a time-dependent neural network $s_{\theta}(\mathbf{x},t)$ to a score function $\nabla_{\mathbf{x}}\log p_{t}{(\mathbf{x})}$. The solution of the forward SDE equation is that:
${L}:=\mathbb{E}_{t}$$\{\lambda(t)\mathbb{E}_{\mathbf{x}_{0}}\mathbb{E}_{q(\mathbf{x}_{t}|\mathbf{x}_{0})}[\|{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t}t)-\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|\mathbf{x}_{0})\|_{2}^{2}]\},$ (16)
@@
where $\mathbf{x}_{0}$ is sampled from distribution $p_{0}$ and $\lambda(t)$ is the positive weighting function. This method circumvents the direct approximation of the computationally infeasible score function by estimating the transition probability that follows a Gaussian distribution during the forward diffusion process [6].
Besides, there are some simplified explanations of reverse-time SDE solvers. Using these techniques, samples can be generated after training using various methods.
  • Euler-Maruyama (EM) Method [147]: Solves the reverse-time SDE through a simple discretization technique, replacing $\mathrm{d}\mathbf{x}$ with $\triangle\mathbf{t}$ and $\mathrm{d}\bar{\mathbf{w}}$ with Gaussian noise $z$.
  • Prediction-Correction (PC) Method: Operates in a sequential manner, alternating between predictor and corrector steps. The predictor can use any numerical solver, such as the EM method, for the reverse-time SDE, while the corrector can be any score-based Markov Chain Monte Carlo (MCMC) method.
  • Probability Flow ODE Method [6]: Reformulates the forward SDE into an ODE that maintains the same marginal probability density $p_{t}$ as the SDE. Sampling by solving this reverse-time ODE is equivalent to solving the time-reversed SDE. There are also some advanced ODE solvers to speed up the process [145, 146].
2.3.3 Conditional Diffusion Models
@@
In the previous sections, we discuss the DDPM and Score SDE from an unconditional view. They generate data without any explicit conditions or guidance. The model learns to produce outputs from the learned distribution of the input data. The general diffusion models are capable of generating data samples not just from an unconditional distribution $p_{0}$, but also from a conditional distribution $p_{0}(\mathbf{x}|c)$ when given a condition $c$. This condition could be class labels or features related to the input data $\mathbf{x}$ [7]. During training, the score network $s_{\theta}(\mathbf{x},t,c)$ takes condition $c$ as an input. Additionally, there are specific sampling algorithms designed for conditional generation, such as label-based conditions [8], label-free conditions [9], and further distillation-based guidance [8], self guidance [150, 171], textual based guidance [149, 172], graph-based guidance [173], physical based guidance [163], task-based guidance [165], etc. These conditional mechanisms are more conducive to the generation of application-specific fields by using the control of other information to generate results [115].
@@
In detail, sampling under labels and classifiers’ conditions involves using gradient guidance at each step, which typically requires an additional classifier with encoder architecture (e.g., U-Net [151] and Transformer [152]) to generate condition gradients for specific labels [8]. These labels are flexible, and can be textual or categorical, binary, or based on extracted features [8, 146, 153, 154, 155, 201, 157, 158, 167, 169]. Correspondingly, sampling under unlabeled conditions relies solely on self-information for guidance [150, 159, 164]. Compared to the high accuracy of the labeled conditional diffusion model, the unlabeled one has advantages in generating innovative and diverse data. Therefore, unlabelled models are more suitable for exploratory and creative application scenarios [160, 161]. Furthermore, there are also more conditional methods being proposed, which use information about the data itself [164], other modalities [153, 263], other representations [163], and other knowledge [168] as conditions to guide the diffusion model for generation. Currently, the condition-based diffusion model is also the most common method in various application scenarios due to its highly specific and controlled outputs [162, 199].
Figure 5: Categorization of improvements with diffusion model and its variants.
2.3.4 Improvements with Diffusion Model and Its Variants

### v5

@@ 2.3. Typical Diffusion Models
Typically, the training process includes two steps: the forward process (diffusion) and the reverse process (denoising). Diffusion models start with a noise distribution, which is gradually altered through a series of steps. In the forward process, the model incrementally adds noise to the original data over multiple steps until the data turns into pure random noise. This process is usually Markovian, meaning that each step depends only on the preceding one. Then the reverse process takes place, involving learning to remove the noise from the data, essentially reversing the forward process. By training the model to remove the noise added during the diffusion process, the model learns to generate samples from the same distribution as the training data. The entire training process involves optimizing the model to denoise effectively. This is typically done using a loss function that encourages the model to produce samples that are close to the true data distribution (Nichol and Dhariwal, 2021).
@@
The currently common frameworks for diffusion models include denoised diffusion probabilistic models (DDPMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020), score-based stochastic differential equations (Score SDEs) (Song et al., 2020b; Song and Ermon, 2019), conditional diffusion models (Rombach et al., 2022; Dhariwal and Nichol, 2021; Ho and Salimans, 2022), etc. Following, we will introduce the subclasses of the diffusion model through theoretical formula derivation.
2.3.1. Denoised Diffusion Probabilistic Models (DDPMs)
@@
(1) $q(\mathbf{x}_{k}|\mathbf{x}_{k-1}):=\mathcal{N}(\mathbf{x}_{k};\sqrt{1-\beta_{k}}\mathbf{x}_{k-1},\beta_{k}\mathbf{I}),$
@@
(2) $q(\mathbf{x}_{k}|\mathbf{x}_{0}):=\prod_{k=1}^{K}{q}(\mathbf{x}_{k}|\mathbf{x}_{k-1}):=\mathcal{N}(\mathbf{x}_{k};\sqrt{\bar{\alpha_{k}}}\mathbf{x}_{0},\sqrt{1-\bar{\alpha_{k}}}\mathbf{I}),$
with $\alpha_{k}:=1-\beta_{k}$ and $\bar{\alpha_{k}}:=\prod_{i=1}^{K}\alpha_{i}$. Therefore, $\mathbf{x}_{k}=\sqrt{\bar{\alpha_{k}}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha_{k}}}\epsilon$ with Gaussian noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. Typically, it is designed $\bar{\alpha_{k}}\approx 0$, s.t., $q(\mathbf{x}_{k}):=\int q(\mathbf{x}_{K}|\mathbf{x}_{0})q(\mathbf{x}_{0})\mathbf{d}\mathbf{x}_{0}\approx\mathcal{N}(\mathbf{x}_{k};\mathbf{0},\mathbf{I})$, i.e., the backward chain can begin with any Gaussian noise. Overall, the forward process gradually injects noise into the data until all structures have disappeared.
@@
(3) ${p}_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k}):=\mathcal{N}(\mathbf{x}_{k-1};\mu_{\theta}(\mathbf{x}_{k},{k}),\sigma_{\theta}(\mathbf{x}_{k},{k})\mathbf{I}),$
@@
(4) $\mathbb{E}\left[-\log p_{\theta}\left(\mathrm{x}_{0}\right)\right]$ $\leq\mathbb{E}_{q}\left[-\log\frac{p_{\theta}\left(\mathrm{x}_{0:K}\right)}{q\left(\mathrm{x}_{1:K}\mid\mathrm{x}_{0}\right)}\right]=\mathbb{E}_{q}\left[-\log p\left(\mathrm{x}_{K}\right)-\sum_{k\geq 1}\log\frac{p_{\theta}\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k}\right)}{q\left(\mathrm{x}_{k}\mid\mathrm{x}_{k-1}\right)}\right]=:L.$
It is equivalent to Kullback–Leibler divergence (KL divergence) format as mentioned in the DDPM paper (Ho et al., 2020) with three parts: the prior loss $L_{K}$, the divergence of the forwarding step and the corresponding reversing step $L_{k-1}$, and the reconstruction loss $L_{0}$:
(5) $L$= $\mathbb{E}_{q}[\underbrace{D_{\mathrm{KL}}\left(q\left(\mathrm{x}_{K}\mid\mathrm{x}_{0}\right)\|p\left(\mathrm{x}_{K}\right)\right)}_{L_{K}}+\sum_{k>1}\underbrace{D_{\mathrm{KL}}\left(q\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k},\mathrm{x}_{0}\right)\|p_{\theta}\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k}\right)\right)}_{L_{k-1}}-\underbrace{\log p_{\theta}\left(\mathrm{x}_{0}\mid\mathrm{x}_{1}\right)}_{L_{0}}].$
@@
(6) $q\left(\mathrm{x}_{k-1}\mid\mathrm{x}_{k},\mathrm{x}_{0}\right)=\mathcal{N}\left(\mathrm{x}_{k-1};\tilde{\boldsymbol{\mu}}_{k}\left(\mathrm{x}_{k},\mathrm{x}_{0}\right),\tilde{\beta}_{k}{I}\right),$
(7) $\tilde{\mu}_{t}(\mathbf{x}_{k},\mathbf{x}_{0}):=\frac{\sqrt{\bar{\alpha}_{k-1}}\beta_{k}}{1-\bar{\alpha}_{k}}\mathbf{x}_{0}+\frac{\sqrt{\alpha_{k}}(1-\bar{\alpha}_{k-1})}{1-\bar{\alpha}_{k}}\mathbf{x}_{k}\hskip 2.84526pt,$
(8) $\tilde{\beta}_{k}:=\frac{1-\bar{\alpha}_{k-1}}{1-\bar{\alpha}_{k}}\beta_{k}.$
@@
(9) $L_{k-1}=\mathbb{E}_{q}\left[\frac{1}{2\sigma_{k}^{2}}\left\|\tilde{\boldsymbol{\mu}}_{k}\left(\mathrm{x}_{k},\mathrm{x}_{0}\right)-\boldsymbol{\mu}_{\theta}\left(\mathrm{x}_{k},k\right)\right\|^{2}\right]+C.$
Ho et al. (Ho et al., 2020) emphasize that, rather than parameterizing the mean $\mu_{\theta}(\mathrm{x}_{k},k)$, predicting the noise vector at each time step in the forward process by parameterizing $\epsilon_{\theta}(\mathrm{x}_{k},k)$ for simplification:
(10) $\mathbb{E}_{k\sim\mathcal{U}(1,K),\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\bigg[{\lambda(k)\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{k},k)\right\|^{2}}\bigg],$
where $\lambda(k)=\frac{{\beta_{k}}^{2}}{2{\sigma_{K}}^{2}\alpha_{k}(1-\bar{\alpha}_{k})}$ is a weight that changes the noise scale and $\epsilon_{\theta}$ is a model for Gaussian noise prediction. After training using the above loss function, $\epsilon_{\theta}$ will be used in the reverse process of ancestor sampling.
@@
(11) $p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k})$ $=\mathcal{N}(\mathbf{x}_{k-1};\mu_{\theta}(\mathbf{x}_{k},k),\sigma_{\theta}^{2}(\mathbf{x}_{k},k)\mathbf{I})\sim\frac{1}{\sqrt{\alpha_{k}}}(\mathbf{x}_{k}-\frac{\beta_{k}}{\sqrt{1-\overline{\alpha_{k}}}}\epsilon_{\theta}(\mathbf{x}_{k},k))+\sigma_{\theta}(\mathbf{x}_{k},k)z,$
@@ 2.3.2. Score SDE Formulation
DDPM achieved a set of discrete steps in the forward processing, so it has some limitations about training designs. Score SDE further generalizes DDPM’s discrete system to a continuous framework based on the stochastical differential equation (Song et al., 2020b). Here we use $T$ instead of the step size $k$ in DDPM.
@@
Forward Process. The corresponding continuous diffusion process can be represented using Itô SDE (Ito et al., 1951), including a mean shift and a Brownian motion (standard Wiener process) as follows:
(12) $\mathrm{d}\mathbf{x}=\mathbf{f}(\mathbf{x},t)\mathrm{d}t+g(t)\mathrm{d}\mathbf{w},t\in[0,T],$
@@
where $\mathbf{f}(\cdot,t)$ represents the drift coefficient for the stochastic process $\mathbf{x}(t)$, and $g(\cdot)$ is the diffusion coefficient linked to the Brownian motion $\mathbf{w}$.
@@
Reverse Process. The new samples can be synthesized from the known prior distribution $p_{T}$ by solving the reverse-time SDE (Anderson, 1982):
(13) $\mathrm{d}\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]\mathrm{d}t+g(t)\mathrm{d}\bar{\mathbf{w}},$
@@
where $\bar{\mathbf{w}}$ is a Brownian motion with reversed time flows (Vincent, 2011). The solution to the reverse-time SDE is approximated by a time-dependent neural network $s_{\theta}(\mathbf{x},t)$ to a score function $\nabla_{\mathbf{x}}\log p_{t}{(\mathbf{x})}$. The solution of the forward SDE equation is that:
(14) ${L}=\mathbb{E}_{t}$ $\{\lambda(t)\mathbb{E}_{\mathbf{x}_{0}}\mathbb{E}_{q(\mathbf{x}_{t}|\mathbf{x}_{0})}[\|{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t}t)-\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|\mathbf{x}_{0})\|_{2}^{2}]\},$
@@
where $\mathbf{x}_{0}$ is sampled from the distribution $p_{0}$ and $\lambda(t)$ is the positive weighting function. This method circumvents the direct approximation of the computationally infeasible score function by estimating the transition probability that follows a Gaussian distribution during the forward diffusion process (Song et al., 2020b).
Besides, there are some simplified explanations of reverse-time SDE solvers. Using these techniques, samples can be generated after training using various methods. Euler-Maruyama (EM) Method (Mao, 2015): Solves the reverse-time SDE through a simple discretization technique, replacing $\mathrm{d}\mathbf{x}$ with $\triangle\mathbf{t}$ and $\mathrm{d}\bar{\mathbf{w}}$ with Gaussian noise $z$. Prediction-Correction (PC) Method: Operates in a sequential manner, alternating between predictor and corrector steps. The predictor can use any numerical solver, such as the EM method, for the reverse-time SDE, while the corrector can be any score-based Markov Chain Monte Carlo (MCMC) method. Probability Flow ODE Method (Song et al., 2020b): Reformulates the forward SDE into an ODE that maintains the same marginal probability density $p_{t}$ as the SDE. Sampling by solving this reverse-time ODE is equivalent to solving the time-reversed SDE. There are also some advanced ODE solvers to speed up the process (Liu et al., 2022c; Lu et al., 2022b).
2.3.3. Conditional Diffusion Models
@@
In the previous sections, we discussed the DDPM and Score SDE from an unconditional view. They generate data without any explicit conditions or guidance. The model learns to produce outputs from the learned distribution of the input data. The general diffusion models are capable of generating data samples not just from an unconditional distribution $p_{0}$, but also from a conditional distribution $p_{0}(\mathbf{x}|c)$ when given a condition $c$. This condition could be class labels or features related to the input data $\mathbf{x}$ (Rombach et al., 2022). During training, the score network $s_{\theta}(\mathbf{x},t,c)$ takes condition $c$ as an input. Additionally, there are specific sampling algorithms designed for conditional generation, such as label-based conditions (Dhariwal and Nichol, 2021), label-free conditions (Ho and Salimans, 2022), and further distillation-based guidance (Dhariwal and Nichol, 2021), self guidance (Epstein et al., 2024; Yang et al., 2022), textual-based guidance (Le et al., 2024; Gong et al., 2022), graph-based guidance (Schneuing et al., 2022), physical-based guidance (Yuan et al., 2023b), task-based guidance (An et al., 2023), etc. These conditional mechanisms are more conducive to the generation of application-specific fields by using the control of other information to generate results (Yang et al., 2023e).
@@
In detail, sampling under labels and classifiers’ conditions involves using gradient guidance at each step, which typically requires an additional classifier with encoder architecture (e.g., U-Net (Ronneberger et al., 2015) and Transformer (Vaswani et al., 2017)) to generate condition gradients for specific labels (Dhariwal and Nichol, 2021). These labels are flexible and can be textual or categorical, binary, or based on extracted features (Dhariwal and Nichol, 2021; Lu et al., 2022b; Nichol et al., 2021; Meng et al., 2023; Hu et al., 2022; Wolleb et al., 2022b; Chen et al., 2023b; Baranchuk et al., 2021; Yang et al., 2024b; Zuo et al., 2023). Correspondingly, sampling under unlabeled conditions relies solely on self-information for guidance (Epstein et al., 2024; Chao et al., 2022; Kollovieh et al., 2024). Compared to the high accuracy of the labeled conditional diffusion model, the unlabeled one has advantages in generating innovative and diverse data. Therefore, unlabelled models are more suitable for exploratory and creative application scenarios (Chung et al., 2022; Choi et al., 2021). Furthermore, there are also more conditional methods being proposed, which use information about the data itself (Kollovieh et al., 2024), other modalities (Nichol et al., 2021; Liu et al., 2023a), other representations (Yuan et al., 2023b), and other knowledge (Shao et al., 2024) as conditions to guide the diffusion model for generation. Currently, the condition-based diffusion model is also the most common method in various application scenarios due to its highly specific and controlled outputs (Bansal et al., 2023; Zhang et al., 2023c).
2.3.4. Improvements with Diffusion Model and Its Variants